Circular No. 8742

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 Q1 (McNAUGHT) R. H. McNaught reports his discovery of a comet (discovery observation

tabulated below) on CCD images obtained with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey, noting a soft image with an apparent coma diameter of ∼ 8′′ in 4′′ seeing; his observations on Aug. 21.73 and 21.78 UT showed the comet to be again slightly diffuse, with deep exposures showing a diffuse condensation and a 20′′-diameter coma. CCD images taken at low altitude by J. Young with the 0.61-m reflector at Table Mountain on Aug. 23.5 show a round, 8′′ coma with a hint of central condensation, with no suggestion of any tail in long exposures.

2006 UT  α_{2000}  δ_{2000}  Mag.
Aug. 20.74130  5^{h}24^{m}12.^{s}81  −25^{o}25′43′′.5  17.8

The initial astrometry and a very preliminary orbit were published on MPEC 2006-Q22. Additional astrometry and revised parabolic orbital elements [T = 2008 June 29.669 TT, q = 2.70330 AU, ω = 345^{o}.338, Ω = 199^{o}.406, i = 58^{o}.994 (equinox 2000.0)] appear on MPEC 2006-Q50.

COMET 73P/SCHWASSMANN-WACHMANN M. L. Sitko, University of Cincinnati and Space Science Institute; R.

W. Russell, D. K. Lynch, and R. Ford, Aerospace Corporation; H. B. Hammel, Space Science Institute; and W. Golisch and P. Sears, Infrared Telescope Facility (IRTF), NASA, report on post-perihelion observations made of comet 73P’s component ‘C’ using the IRTF (+ BASS) on Aug. 5, 7, and 8 UT. The comet exhibited a continuum between 3 and 13 μm, on top of which a silicate emission band from 8.5 to 12.2 μm was observed. The flux observed on all three days was approximately the same, to within 10 percent. An underlying blackbody, normalized to the continuum fluxes of the averaged spectra at 8.1 and 12.5 μm, yielded a mean grain temperature of 272 K (estimated uncertainty ± 5 K). The derived temperature was 10– 11 percent higher than that of an equivalent blackbody at the heliocentric distance of the comet. The feature-to-continuum ratio in the silicate band was 1.19, smaller than that previously reported for this fragment from preperihelion observations (IAUC 8701 and 8717). The measured flux between 10.0 and 11.0 μm, using the 3′′.4 circular entrance aperture of BASS, was 0.28 ± 0.03 Jy (equivalent magnitude m_{N} = 5.4 ± 0.1).

2006 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green